# Understanding Hallucinations in Diffusion Models through Mode Interpolation

### v1

@@ Abstract
Colloquially speaking, image generation models based upon diffusion processes are frequently said to exhibit “hallucinations,” samples that could never occur in the training data. But where do such hallucinations come from? In this paper, we study a particular failure mode in diffusion models, which we term mode interpolation. Specifically, we find that diffusion models smoothly “interpolate” between nearby data modes in the training set, to generate samples that are completely outside the support of the original training distribution; this phenomenon leads diffusion models to generate artifacts that never existed in real data (i.e., hallucinations). We systematically study the reasons for, and the manifestation of this phenomenon. Through experiments on 1D and 2D Gaussians, we show how a discontinuous loss landscape in the diffusion model’s decoder leads to a region where any smooth approximation will cause such hallucinations. Through experiments on artificial datasets with various shapes, we show how hallucination leads to the generation of combinations of shapes that never existed. Finally, we show that diffusion models in fact know when they go out of support and hallucinate. This is captured by the high variance in the trajectory of the generated sample towards the final few backward sampling process. Using a simple metric to capture this variance, we can remove over 95% of hallucinations at generation time while retaining 96% of in-support samples. We conclude our exploration by showing the implications of such hallucination (and its removal) on the collapse (and stabilization) of recursive training on synthetic data with experiments on MNIST and 2D Gaussians dataset. We release our code at https://github.com/locuslab/diffusion-model-hallucination.
@@ 1 Introduction
In this work, we formalize and study a particular failure mode of diffusion models that we call hallucination—a phenomenon where diffusion models generate samples that lie completely out of the support of the training distribution of the model. We observe hallucinations even in large generative models like StableDiffusion [28] in the form of hands with extra (or missing) fingers or limbs. We begin our investigation with a surprising observation that an unconditional diffusion model trained on a distribution of simple shapes, generates images with combinations of shapes (or artifacts) that never existed in the original training distribution (Figure 1). While a lot of research in generative models has studied the phenomenon of ‘mode collapse’ [43] which leads to a loss in diversity in the sampled distribution, in our work we explain hallucinations through the lens of a phenomenon we call ‘mode interpolation’, which has not been studied in past work to the best of our knowledge.
@@
To understand the cause of these hallucinations and their relationship to mode interpolation, we construct simplified 1-d and 2-d mixture of Gaussian setups and train diffusion models on them (§ 4). We observe that when the true data distribution occurs in disjoint modes, diffusion models are unable to model a true approximation of the underlying distribution. This is because there exist ‘step functions’ between different modes, but the score function learned by the DDPM is a smooth approximation of the same, leading to interpolation between the nearest modes, even when these interpolated values are entirely absent from the training data. Our investigation on the occurrence of mode interpolation leads us to the observation that hallucinated samples usually have very high variance towards the end of their trajectory during the reverse diffusion process. Based on this observation, we use the trajectory variance during sampling as a metric to detect hallucinations (§ 5), and show that diffusion models usually ‘know’ when they hallucinate, allowing detection with sensitivity and specificity $>0.92$ in our experiments.
@@
Finally, we study the implications of this phenomenon in recursive generative model retraining where we train generative models on their own output (§ 6). Recently, recursive training and its downsides in model collapse have garnered a lot of attention in both language and diffusion modeling literature [1, 2, 7, 4]. We observe that the proposed detection mechanism is able to mitigate the model collapse during recursive training on 2D Grid of Gaussians, Shapes and MNIST dataset.
@@
Figure 1: Hallucinations in Diffusion Models: Original Dataset (Left) & Generated Dataset (Right). The original dataset consists of 64x64 images divided into three columns, each containing a triangle, square, or pentagon with a 0.5 probability of the shape being present. Each shape appears at most once per image. The generated dataset created using an unconditional DDPM includes some samples (hallucinations) with multiple occurrences of the same shape that is unseen in the original dataset.
@@ 4 Understanding Mode Interpolation and Hallucination
The rate of mode interpolation depends primarily on three factors: (i) Number of training data points, (ii) variance of (and distance between) the distributions, and (iii) the number of sampling timesteps ($T$). As the number of training samples increases, we observe that the proportion of interpolated samples decreases. In this setup, the variance of $p(x)$ not only depends on $\sigma$ but also the distance between the modes i.e, $|\mu_{1}-\mu_{2}|$ and $|\mu_{2}-\mu_{3}|$. We run another experiment with $\mu_{1}=1,\mu_{2}=2\text{ and }\mu_{3}=4$. In this case, we observe that the frequency of samples between $\mu_{2}$ and $\mu_{3}$ is much lower than $\mu_{1}$ and $\mu_{2}$. The number of interpolated samples also decreases as the distance from the modes increases. The frequency of interpolated samples is also proportional to the number of timesteps $T$. Additional experiments with different numbers of Gaussians are presented in Appendix C.
@@ 5.2 Metric for detecting hallucination
Based on the above observation about high variance in predicted values of $x_{0}$ in the reverse diffusion process, we use the same observation as a metric to distinguish hallucinated and non-hallucinated (in-support) samples. Mathematically, the metric can be defined as follows:
$\texttt{Hal}(x)=\frac{1}{T}\sum_{i=0}^{T}(\hat{x_{0}}^{(t)}-\overline{{\hat{x_{0}}}^{(t)}})^{2}$ (4)
where $\hat{x_{0}}^{(t)}$ represents the predicted values of the final image at different time steps $(t)$, and $\overline{{\hat{x_{0}}}^{(t)}}$ is the mean of these predictions over the same time steps. We now utilize this metric to analyze the histogram values of each sample from the three experimental setups studied thus far.
@@
Finally, we conclude our investigation on synthetic datasets with experiments on the 2D Gaussian dataset. Similar to the 1D Gaussian setup, we once again measure the prediction variance of the last 20 steps of the reverse diffusion process. We compute the variance per dimension and then take the mean across dimensions to . With this metric, we can filter out 96% of the hallucinated samples while retaining 95% of the in-support samples.
@@ 7 Discussion
In this work, we performed an in-depth study to formulate and understand hallucination in diffusion models, focusing on the phenomenon of mode interpolation. We analyzed this phenomenon in three different settings: 1D Gaussian, 2D Grid of Gaussians, and the shapes dataset, and saw how diffusion models learn smoothed approximations of disjoint score functions, leading to mode interpolation. Based on our analysis, we developed a metric to identify hallucinated samples effectively and explored the implications of hallucination in the context of recursive generative model training. Past works [30] have shown that rare concepts/classes in large diffusion models like StableDiffusion [28] are poorly modeled. This is evident from the distortion of hands commonly observed in samples generated by these models. We suspect that this might be a case of mode interpolation. The occurrence of 6-8 fingers is potentially analogous to the occurrence of 2 squares in the toy experiment. We hope our work inspires future research in understanding and mitigating hallucination in diffusion models.
@@ A.3 MNIST
Code: We provide code to run all experiments in the supplementary material.
@@ Appendix B Limitations and Broader Impact
Most of our evidence for mode interpolation comes from the 1D/2D synthetic toy setups. We do not clearly demonstrate what mode interpolation and hallucination look like in real-world natural images. This is a challenging problem because natural images have a high-dimensional, complex distribution.
@@ Appendix C Additional Experiments and Figures
    We also experiment with mixture of 2 Gaussians in Figure  11 and 4 Gaussians in Figure  12.

### v2

@@ Abstract
Colloquially speaking, image generation models based upon diffusion processes are frequently said to exhibit “hallucinations”—samples that could never occur in the training data. But where do such hallucinations come from? In this paper, we study a particular failure mode in diffusion models, which we term mode interpolation. Specifically, we find that diffusion models smoothly “interpolate” between nearby data modes in the training set to generate samples that are completely outside the support of the original training distribution; this phenomenon leads diffusion models to generate artifacts that never existed in real data (i.e., hallucinations). We systematically study the reasons for, and the manifestation of this phenomenon. Through experiments on 1D and 2D Gaussians, we show how a discontinuous loss landscape in the diffusion model’s decoder leads to a region where any smooth approximation will cause such hallucinations. Through experiments on artificial datasets with various shapes, we show how hallucination leads to the generation of combinations of shapes that never existed. We extend the validity of mode interpolation in real-world datasets by explaining the unexpected generation of images with additional or missing fingers similar to those produced by popular text-to-image generative models. Finally, we show that diffusion models in fact know when they go out of support and hallucinate. This is captured by the high variance in the trajectory of the generated sample towards the final few backward sampling steps. Using a simple metric to capture this variance, we can remove over 95% of hallucinations at generation time while retaining 96% of in-support samples in the synthetic datasets. We conclude our exploration by showing the implications of such hallucination (and its removal) on the collapse (and stabilization) of recursive training on synthetic data with experiments on MNIST and a 2D Gaussians dataset. We release our code at https://github.com/locuslab/diffusion-model-hallucination.
@@ 1 Introduction
In this work, we formalize and study a particular failure mode of diffusion models that we call hallucination—a phenomenon where diffusion models generate samples that lie completely out of the support of the training distribution of the model. As a contemporary example, hallucinations manifest in large generative models like StableDiffusion [32] in the form of hands with extra (or missing) fingers or limbs. We begin our investigation with a surprising observation that an unconditional diffusion model trained on a distribution of simple shapes, generates images with combinations of shapes (or artifacts) that never existed in the original training distribution (Figure 1). While extensive research on generative models has focused on the phenomenon of ‘mode collapse’ [47], which leads to a loss of diversity in the sampled distribution, such studies often overlook the complex nature of real data which typically comprise multiple distinct modes on a complex data manifold, and the effects of their mutual interactions are thus neglected. In our work, we explain hallucinations by introducing a novel phenomenon we term ‘mode interpolation’ that considers this mutual interaction.
@@
To understand the cause of these hallucinations and their relationship to mode interpolation, we construct simplified 1-d and 2-d mixture of Gaussian setups and train diffusion models on them (§ 4). We observe that when the true data distribution occurs in disjoint modes, diffusion models are unable to model a true approximation of the underlying distribution. This is because there exist ‘step functions’ between different modes, but the score function learned by the DDPM is a smooth approximation of the same, leading to interpolation between the nearest modes, even when these interpolated values are entirely absent from the training data. Moreover, we observation that hallucinated samples usually have very high variance towards the end of their trajectory during the reverse diffusion process. Based on this observation, we use the trajectory variance during sampling as a metric to detect hallucinations (§ 5), and show that diffusion models usually ‘know’ when they hallucinate, allowing detection with sensitivity and specificity $>0.92$ in our experiments.
@@
We explore mode interpolation as a potential explanation for the common failure of large-scale generative models, to accurately generate human hands. To demonstrate this concretely, we trained a diffusion model on a dataset of high-quality hand images and observed that it generated hands with additional fingers. We then applied our proposed metric to effectively detect these hallucinated generations. Finally, we study the implications of this phenomenon in recursive generative model retraining where we train generative models on their own output (§ 6). Recently, recursive training and its downsides in model collapse have garnered a lot of attention in both language and diffusion modeling literature [2, 3, 8, 5]. We observe that the proposed detection mechanism is able to mitigate the model collapse during recursive training on 2D Grid of Gaussians, Shapes and MNIST dataset.
@@
Figure 1: Hallucinations in Diffusion Models: Original Dataset (Left) & Generated Dataset (Right). (Top) The original dataset consists of 64x64 images divided into three columns, each containing a triangle, square, or pentagon with a 0.5 probability of the shape being present. Each shape appears at most once per image. The generated dataset created using an unconditional DDPM includes some samples (hallucinations) with multiple occurrences of the same shape that is unseen in the original dataset. (Bottom) We also train a ADM [28] on a dataset of high-quality images of human hands and show that the diffusion model generates hallucinated images of hands with additional fingers.
@@ 4 Understanding Mode Interpolation and Hallucination
The rate of mode interpolation depends primarily on three factors: (i) Number of training data points, (ii) variance of (and distance between) the distributions, and (iii) the number of sampling timesteps ($T$). As the number of training samples increases, we observe that the proportion of interpolated samples decreases. In this setup, the variance of $p(x)$ not only depends on $\sigma$ but also the distance between the modes i.e, $|\mu_{1}-\mu_{2}|$ and $|\mu_{2}-\mu_{3}|$. We run another experiment with $\mu_{1}=1,\mu_{2}=2\text{ and }\mu_{3}=4$. In this case, we observe that the frequency of samples between $\mu_{2}$ and $\mu_{3}$ is much lower than $\mu_{1}$ and $\mu_{2}$. The number of interpolated samples also decreases as the distance from the modes increases. The frequency of interpolated samples is also inversely proportional to the number of timesteps $T$. Additional experiments with different numbers of Gaussians are presented in Appendix C.
@@ 4.3 What causes mode interpolation?
4.4 Simple Shapes
We now discuss the mode interpolation in the Simple Shapes dataset. In this context, the interpolation is not happening in the output space, but rather in the representation space. To investigate this, we performed a t-SNE visualization of the outputs from the bottleneck layer of the U-Net used in the Simple Shapes experiment, as shown in Figure  10 . Regions 1 and 3 in the representation space semantically correspond to the images where squares are at the top and bottom of the image respectively. At inference time, we can see a clear emergence of region 2 which is between regions 1 and 3 (interpolated), and contains two squares (hallucinations) at the top and bottom of the image. This experiment concretely confirms that interpolation happens in representation space.
Figure 5: Hands Dataset. We train a ADM on the Hands dataset with 5000 images (first column) and show that the generated samples (second column) consists of hallucinated samples (additional/missing fingers). We then apply our proposed metric to detect these hallucinated samples (third column).
4.5 Mode Interpolation in Real World datasets: Hands
We sought to demonstrate the occurrence of mode interpolation in a real-world setting. A well-documented challenge with popular text-to-image generative models is their difficulty in accurately generating human hands [27]. Despite extensive research in modern diffusion models, there is no conclusive explanation for the missing/additional fingers generated by these models. One hypothesis attributes this difficulty to the anatomical complexity of human hands, which involve numerous joints, fingers, and diverse poses. Another hypothesis suggests that, although large datasets contain many images of hands, these hands are often partially obscured (e.g., when a person is holding a cup) and occupy only a small region of the overall image.
To investigate this further, we trained a diffusion model on a datasets with high-quality images of human hands. The Hands dataset [1] consists of high resolution images of hands from 190 subjects of various ages. Each subject’s right and left hands were photographed while opening and closing fingers against a uniform white background. We sample 5000 images from the Hands dataset and train an ADM [28] model on this dataset. We resize the images to 128x128 and use the same hyperparameters as that of the FFHQ dataset [17]. We mention all the hyperparameters in the Appendix A. We observe images with additional and missing fingers in the generated samples as seen in Figure  5. This is a pretty surprising result as it is non-trivial to assume that diffusion model generates images with additional fingers. Despite the potential for various failure modes, such as blurred hand images, these issues were not observed in our results. In some ways, the occurrence of 6-8 fingers is analogous to the occurrence of 2 squares in the Simple Shapes dataset. Thus, the presence of additional fingers in these images (i.e hallucinated images) generated by the diffusion model demonstrates the phenomenon of mode interpolation in real-world datasets. More example are shown in Fig.  20 &  21.
@@ 5.2 Metric for detecting hallucination
Based on the above observation about high variance in predicted values of $x_{0}$ in the reverse diffusion process, we use the same observation as a metric to distinguish hallucinated and non-hallucinated (in-support) samples. The intuition behind the metric is to capture the variance in the trajectory of $\hat{x_{0}}$. Let $T_{1}$ be the starting timestep and $T_{2}$ be the end timestep. Mathematically, the metric can be defined as follows:
$\texttt{Hal}(x)=\frac{1}{|T_{2}-T_{1}|}\sum_{i=T_{1}}^{T_{2}}\left(\hat{x_{0}}^{(i)}-\overline{{\hat{x_{0}}}^{(t)}}\right)^{2}$ (4)
where $\hat{x_{0}}^{(t)}$ represents the predicted values of the final image at different time steps $(t)$, and $\overline{{\hat{x_{0}}}^{(t)}}$ is the mean of these predictions over the same time steps. We now utilize this metric to analyze the histogram values of each sample from the three experimental setups studied thus far. This metric can be implemented in two ways. One approach is to store $\hat{x_{0}}$ during the reverse diffusion process and then compute the variance. Alternatively, we explore a method where forward diffusion is performed for $k$ steps between $T_{1}$ and $T_{2}$, predicting $\hat{x_{0}}$ at each step, and then computing the variance.
@@
Next, we discuss our investigation on synthetic datasets with experiments on the 2D Gaussian dataset. Similar to the 1D Gaussian setup, we once again measure the prediction variance of the last 20 steps of the reverse diffusion process. We compute the variance per dimension and then take the mean across dimensions to . With this metric, we can filter out 96% of the hallucinated samples while retaining 95% of the in-support samples.
Hands
Finally, we conclude our investigation with experiments on the Hands dataset. To analyze the effectiveness of the proposed metric, we manually label  130 images from the generated samples as hallucinated vs. in-support. This includes 88 images with 5 fingers and 40 images with missing/ additional fingers i.e. hallucinated samples. The histogram (in Figure  5) shows that the proposed metric can indeed detect these hallucinations to a reasonable degree. In our experiments, we observe that we can eliminate  80% of the hallucinated samples while retaining  81% of the in-support samples. The trajectories of the hallucinated and in-support samples are shown in Figures  22 and  23, respectively. A higher variance in the trajectory of $\hat{x_{0}}$ is clearly observed in the hallucinated samples compared to the in-support samples. We note that the detection is a hard problem and the fact that the method transfers to the real world is proof of the relationship between mode interpolation and hallucination in real-world data.
@@ 7 Discussion
In this work, we performed an in-depth study to formulate and understand hallucination in diffusion models, focusing on the phenomenon of mode interpolation. We analyzed this phenomenon in four different settings: 1D Gaussian, 2D Grid of Gaussians, Shapes and Hands datasets, and saw how diffusion models learn smoothed approximations of disjoint score functions, leading to mode interpolation. Based on our analysis, we developed a metric to identify hallucinated samples effectively and explored the implications of hallucination in the context of recursive generative model training. This study is the first to propose mode interpolation as a potential hypothesis for explaining the generation of additional fingers in large-scale generative models. We hope that future research will build upon this hypothesis and develop methods to mitigate these issues in generative models. We hope our work inspires future research in understanding and mitigating hallucination in diffusion models.
@@ A.3 MNIST
A.4 Hands
We trained for a total of 200k iterations with batch size 16 and a learning rate of 1e-4. The diffusion process was trained with 1000 timesteps ($T=1000$) with a cosine noise schedule. The U-Net comprised 256 channels, with an attention mechanism incorporating 64 channels per head and 3 residual blocks. For sampling, we use 500 timesteps with respacing. We base our implementation and hyperparameters on the official DDPM-IP [29] repository.
@@ Appendix B Limitations and Broader Impact
In current text-to-image generative models, the poorly modeled “hands” are a clear giveaway in identification of AI generated images. The detection of such AI-generated content would be made much more difficult if these hallucinations were identified and removed from the generated images. While our work builds an understanding of hallucinations, and allows us to also detect them, we believe that future generations of models would have become more robust to such hallucinations by virtue of training on more data independent of this work.
Concerning the limitations of the proposed hallucination metric, the selection of the right timesteps is key to be able to detect hallucinations. More analysis on what region of trajectory leads to hallucinations would be useful across various schedules and sampling algorithms. We believe these are great areas for future work to explore. Additional explorations of mode interpolation and hallucinations in real-world datasets would be useful to the community.
@@ Appendix C Additional Experiments and Figures
We also study Variational Diffusion Models (VDM) [19] to verify the generality of our findings. Our results show that the over-smoothed score function phenomenon persists in VDM, supporting the hypothesis that this issue is not specific to DDPM. We train a simple VDM on the 2D Gaussian with 10k samples. We follow the setup and hyperparameters in the official implementation  [^3] https://github.com/google-research/vdm. We train both continuous and discrete variants of VDM on the 2D Gaussian dataset. The main observation is that VDM mitigates the hallucinations significantly especially with more training data but the phenomenon of mode interpolation still exists. In this figure, we also show the impact of the number of sampling steps on the count of hallucinations. We clearly see that increasing the number of sampling steps reduces the number of hallucinated samples. This can be clearly observed in Figure  11 (first two columns) where the count of hallucinations decreases mode interpolation.
@@
    We also experiment with mixture of 2 Gaussians in Figure  14 and 4 Gaussians in Figure  15.
@@
Figure 10: Interpolation in Representation Space. We analyze the bottleneck of the U-Net to demonstrate mode interpolation in the Shapes dataset. We clearly see that Region 2 (which consists of 2 squares) is interpolating between Region 1 (one square in the bottom half) and Region 3 (one square in the top half).
Figure 11: Variational Diffusion Model. We train a Variational Diffusion Model (VDM) on the 2D Gaussian Data with 10k samples (first three columns). $T$ denotes the timesteps during training and $T^{\prime}$ denotes the sampling timesteps. $T=\infty$ refers to the continuous time variant. The fourth column shows a DDPM trained on a 2D Gaussian with imbalanced modes. The boxes indicate the modes with less data. The last column shows result of sampling from the true score function.
@@
Figure 20: Hallucinated Images of Hands generated by the diffusion model.
Figure 21: In-Support Images of Hands generated by the diffusion model.
Figure 22: $\hat{x_{0}}$ trajectory for Hallucinated Sample (with 250 timesteps). We observe high variance/instability during the steps $t=600$ to $t=900$.
Figure 23: $\hat{x_{0}}$ trajectory for In-Support Sample (with 250 timesteps). We do not observe high variance/instability during the steps $t=600$ to $t=900$.